# Hardware-Compatible Single-Shot Feasible-Space Heuristics for Solving the Quadratic Assignment Problem

### v1

@@ 2.1.1 Solution Space
Number of Facilities
and Locations ($n$)
Number of
Binary Variables ($n^{2}$)
Solution Space:
Native ($n!$)
Solution Space:
Binary ($2^{n^{2}}$)
@@
In practice, however, it is often the case that significantly more than three infeasible configurations are required before a new feasible solution is found, as now all feasible solutions exist as local optima, surrounded by high-energy (i.e., penalized by $\lambda$) infeasible solutions. Figure 2.1 illustrates this phenomenon. In Figure 2.1, the tabu-based QUBO heuristic is able to see only 15 feasible solutions, indicated by the line segments, over 1000 iterations. Annealing-based QUBO solvers often have even worse performance, particularly as the temperature decreases and they are forced to accept only moves that result in an improvement.
Figure 2.1: Hamming distance between the solutions observed by three heuristic solvers and the optimal solution for a QAPLIB instance with 15 facilities and locations (225 binary variables in total). The search heuristics used are native space tabu search, a random generator of feasible solutions, and a QUBO space tabu search.
@@ 3.2.2 Feasible-Space Updates for Semi-symmetric Instances
Let $\hat{D}=\frac{1}{2}D+\frac{1}{2}D^{T}$. Since $F$ is symmetric, each $n\times n$ block submatrix within $\hat{D}\otimes F$ is symmetric:
@@ 4.2.2 Updates Executed Using Matrices
Let the members in the set $\mathcal{N}(x)$ be initialized to follow the lexicographic order. The neighbours in Tables 4.1 and 4.2 show examples of the lexicographic ordering. Let the matrices $S_{12},S_{34}\in\{0,1\}^{\binom{n}{2}\times n^{2}}$, where the supports of the $i$-th rows are
@@
4:     $c_{1}\leftarrow n\cdot c+(\bar{z}_{2}\bmod n)$
5:     $c_{2}\leftarrow n\cdot c+(\bar{z}_{1}\bmod n)$
@@ 4.3 Vectorized Gradient Evaluation of the Full Neighbourhood for
Given a feasible $x$ and its neighbour $y$, recall, from Proposition 3.1, that semi-symmetric instances allow for the simplification
@@ 5 Numerical Experiments
We let $\mathcal{T}$ represent the set of trials, or repeats, that is, the number of times an instance is called by Algorithm 4.5. The best optimality gap is then quantifiable by the minimum of the relative optimality gap:
@@ 5.1 Heuristics
We evaluate four heuristics that can be employed in the Choose_Move function (line 9 in Algorithm 4.5). The first heuristic, that we label “Greedy”, is the greedy approach. Given a set of candidates $\mathcal{N}(x)$ and their corresponding value of $\Theta$, we select the one that results in the smallest value, that is,
@@
The second heuristic is a pseudo-greedy approach, that we label “Top10”. This heuristic randomly chooses one neighbour from the set of neighbours that yields the top 10 best elements in $\Theta$. Let $\widehat{\Theta}=\texttt{List}_{\texttt{sorted}}(\Theta)$, with $\Theta$ computed using $\mathcal{N}(x)$. Let $\pi\in\Pi\left(\binom{n}{2}\right)$ be the permutation that yields the ordered vector $\widehat{\Theta}=\Theta(\pi)$. We then select a neighbour at random:
@@
Third, we use a random walk heuristic inspired by the well-known WalkSAT heuristic [28] in the area of satisfiability problems. Given a probability $p\in(0,1)$, this heuristic chooses a neighbour based on the following rule:
@@
Finally, we use the tabu search heuristic [31], employing the label “Tabu”. In brief, a tabu search constructs a so-called tabu list, which records the feasible points visited up to a given limit. Once $\widehat{\Theta}=\texttt{List}_{\texttt{sorted}}(\Theta)$ is computed, this heuristic sequentially iterates over the neighbours within the full neighbourhood. If a neighbour $i$ results in the value of $f(x)+\widehat{\Theta}_{i}$ being better than the best objective value found so far, we select that neighbour for the next iteration. Otherwise, we choose the incumbent neighbour, provided it is not in the tabu list. Once a neighbour is selected, it is appended to the tabu list, and if the tabu list is at its size limit, the oldest item in the list is removed.
5.1.1 Typical Behaviour of the Greedy Search Heuristic
Given a feasible binary solution $x$, the greedy heuristic iteratively selects the neighbour within $\mathcal{N}(x)$ that yields the maximum immediate reduction in the objective function. While this strategy ensures the best possible improvement at each step within the incumbent full neighbourhood, if often leads to an oscillatory pattern, where the search alternates between two permutation assignments. As a result, the algorithm fails to make further progress towards a global optimal solution. We observe this pattern for five of the selected QAPLIB instances in Figure 5.1. We make observations for both settings of $\Theta$, namely, $\Delta$ and $\tilde{\Delta}$.
Figure 5.1: Convergence of the Greedy heuristic on five QAPLIB instances over 50 iterations of full-neighbourhood evaluations.
From the seventh iteration of QAPLIB instance “chr12a”, when the exact gradient is used, the greedy search alternates between the permutation assignments $[5,6,11,1,9,0,7,10,2,8,4,3]\in\Pi_{12}$ and $[5,6,11,1,9,0,7,8,2,10,4,3]\in\Pi_{12}$, yielding objective values of 14,464 and 14,430, respectively. Similarly, when the approximate gradient is used, the greedy search alternate between $[2,6,9,5,11,8,10,3,0,1,4,7]\in\Pi_{12}$ and $[2,6,9,5,11,8,10,3,1,0,4,7]\in\Pi_{12}$, starting from the first iteration, yielding objective values $50266$ and $48284$, respectively. Such behaviour is a commonly observed pattern and it indicates that the greedy search heuristic struggles to escape local minima. Consequently, we do not consider greedy search for further analysis.
@@ 5.2 Comparison of Exact and Approximate Gradients
In this section, we discuss the results of our experiments using the Top10 heuristic (5.2) for two settings of $\Theta$. The number of repeats was set to $100$, that is, $|\mathcal{T}|=100$. The Intel Xeon Gold 6548Y+ processor with 128 cores and 504 GB of memory was used for our experiments. Our aim is to assess the impact of omitting the error correction term $\mathcal{E}$ presented in Proposition 4.4. The primary motivation for omitting this step stems from its computational cost. Computing $\mathcal{E}$ involves a matrix–matrix multiplication $S_{34}Q$, which requires more-substantial processing resources than a matrix–vector multiplication. These operations significantly increase the overall computation time, especially as the problem size grows. The second reason is that when it comes to choosing a neighbour in a given full neighbourhood, the relative ordering of the values in $\Theta$ is important, not the values themselves. For example, given $\Theta^{1}=(-70,-40,-100,3,5)$ and $\Theta^{2}=(-80,-50,-90,3,5)$ representing two distinct gradients, we select the neighbour associated with the smallest value, which is the third element of each vector. Finally, IMC hardware devices typically have limited bit-precision capabilities. As a result, quantization of the nonzero elements in $Q$ becomes necessary. This introduces another potential source of error due to the loss of precision. Consequently, it is crucial to design algorithms that are resilient to multiple sources of error. Despite these challenges, incorporating the error correction step may still be feasible in certain scenarios, particularly when matrix–matrix multiplications can be efficiently performed.
@@
Figure 5.2(a) shows a comparison of the relative optimality gaps using the approximate gradient $\tilde{\Delta}$ and exact gradient $\Delta$ observed over $10^{5}$ iterations, that is, $\left\{{\operatorname*{\text{rel\_gap}}}^{t}(10^{5}):t\in\mathcal{T}\right\}$. Figure 5.2(b) shows the best observed optimality gaps, $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{4})$ and $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{5})$, for both $\tilde{\Delta}$ and $\Delta$. Both approaches yield optimality gaps that are smaller than $1\%$, demonstrating their effectiveness in reaching optimal or near-optimal solutions. This observation supports the use of the approximate gradient as a viable alternative to the exact gradient. However, when comparing the relative runtime of computing the exact versus approximate gradients, the advantage of using the approximate gradient becomes evident. Figure 5.3 shows the relative runtime, computed as the ratio of the time taken for $\Delta$ divided by the time taken for $\tilde{\Delta}$ for each instance, with the instances sorted by size. The plot shows that using the exact gradient requires significantly more computational resources and takes considerably longer than the approximate gradient. Notably, the relative time increase becomes more pronounced as the instance size grows.
Figure 5.3: Ratio of the average runtime using the exact gradient versus the approximate gradient across $100$ trials.
@@
Figure 5.4 shows rel_error over the first $1000$ iterations for instances formulated by Skorin-Kapov (“sko”), Li and Pardalos (“lipa”), and Taillard (“tai”). We observe a consistent trend across all three classes, indicating that the contribution of $\mathcal{E}_{i}$ relative to $\Delta_{i}$ diminishes as the problem size increases, hence, the impact of omitting the error terms becomes negligible in large-scale instances.
@@
5.3 Comparison of Different Heuristics
In this section, we assess the effectiveness of the approximate gradient computation using the heuristics Top10, WalkQAP, and Tabu. The Intel Xeon E7-8890 v3 running at 2.50 GHz with 144 cores and 2.95 TB of memory was used for the experiment. For the WalkQAP heuristic, we set the probability parameter $p=0.95$. For the Tabu heuristic, we set the maximum length of the tabu list to $20$, with the oldest entry being removed when a new neighbour is added whenever the list exceeds this limit. Figure 5.5 presents the relative optimality gaps $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{5})$ for selected instances from the QAPLIB. The results indicate that the Top10 and WalkQAP heuristics achieve better performance compared to the Tabu heuristic.
[…]

### v2

@@ 2.1.1 Solution Space
Number of Facilities Number of Solution Space: Solution Space:
and Locations ($n$) Binary Variables ($n^{2}$) Native ($n!$) Binary ($2^{n^{2}}$)
@@
In practice, however, it is often the case that significantly more than three infeasible configurations are required before a new feasible solution is found, as now all feasible solutions exist as local optima, surrounded by high-energy (i.e., penalized by $\lambda$) infeasible solutions. Figure 2.2 illustrates this phenomenon. In Figure 2.1, the tabu-based QUBO heuristic is able to see only fewer than 30 feasible solutions, indicated by the red markers, over 1000 iterations. Annealing-based QUBO solvers often have even worse performance, particularly as the temperature decreases and they are forced to accept only moves that result in an improvement.
Figure 2.1: Relative optimality gaps between objective values observed by three heuristic solvers and optimal value for the instance “chr12a” of QAPLIB, each having 12 facilities and locations (144 binary variables in total). The search heuristics used are native-space tabu search, a random feasible-solution generator, and a QUBO-space tabu search.
@@ 3.2.2 Feasible-Space Updates for Semi-symmetric Instances
Let $\hat{D}=\frac{1}{2}D+\frac{1}{2}D^{T}$. Since $F$ is symmetric, each $(i,j)$-th $n$-by-$n$ block submatrix within $\hat{D}\otimes F$ is symmetric:
@@ 4.2.2 Updates Executed Using Matrices
Let the members in the set $\mathcal{N}(x)$ be initialized to follow the lexicographic order. The neighbours $N_{0},N_{1},\ldots,N_{9}$ in Tables 4.1 and 4.2 show examples of the lexicographic ordering, which is determined by the first two elements $z_{1}$ and $z_{2}$. Let the matrices $S_{12},S_{34}\in\{0,1\}^{\binom{n}{2}\times n^{2}}$, where the supports of the $i$-th rows are
@@
4:     $c_{1}\leftarrow n\cdot c+(\bar{z}_{1}\bmod n)$
5:     $c_{2}\leftarrow n\cdot c+(\bar{z}_{2}\bmod n)$
@@ 4.3 Vectorized Gradient Evaluation of the Full Neighbourhood for
Given a feasible solution $x$ and its neighbour $y$, recall, from Proposition 3.1, that semi-symmetric instances allow for the simplification
@@ 5 Numerical Experiments
We let $\mathcal{T}$ represent the set of trials, or repeats, that is, the number of times an instance is called by Algorithm 4.5, and ${\operatorname*{\text{rel\_gap}}}^{t}$ be the relative gap obtained at the $t$-th trial. The best optimality gap is then quantifiable by the minimum of the relative optimality gap:
@@ 5.1 Heuristics
We evaluate five heuristics that can be employed in the Choose_Move function (line 9 in Algorithm 4.5). The first heuristic, labelled “Greedy”, is the greedy approach. Given a set of candidates $\mathcal{N}(x)$ and their corresponding value of $\Theta$, we select the one that results in the smallest value, that is,
@@
The second heuristic is a pseudo-greedy approach, labelled “Top10”. This heuristic randomly chooses one neighbour from the set of neighbours that yields the top 10 best elements in $\Theta$. Let $\widehat{\Theta}=\texttt{List}_{\texttt{sorted}}(\Theta)$, with $\Theta$ computed using $\mathcal{N}(x)$. Let $\pi\in\Pi\left(\binom{n}{2}\right)$ be the permutation that yields the ordered vector $\widehat{\Theta}=\Theta(\pi)$. We then select a neighbour at random:
@@
The third heuristic we use is a random walk heuristic inspired by the well-known WalkSAT heuristic [28] in the area of satisfiability problems. Given a probability $p\in(0,1)$, this heuristic chooses a neighbour based on the following rule:
@@
The fourth heuristic is the tabu search heuristic [31], labelled “Tabu”. In brief, a tabu search constructs a so-called tabu list, which records the feasible points visited up to a given limit. Once $\widehat{\Theta}=\texttt{List}_{\texttt{sorted}}(\Theta)$ is computed, this heuristic sequentially iterates over the neighbours within the full neighbourhood. If a neighbour $i$ results in the value of $f(x)+\widehat{\Theta}_{i}$ being better than the best objective value found so far, we select that neighbour for the next iteration. Otherwise, we choose the incumbent neighbour, provided it is not in the tabu list. Once a neighbour is selected, it is appended to the tabu list, and if the tabu list is at its size limit, the oldest item in the list is removed.
The final heuristic we use is simulated annealing, labelled “SA”. The initial temperature $T_{\text{high}}$ is estimated as $\delta_{h}/\log(p_{\text{initial}})$, where $\delta_{h}$ is the $\gamma_{\text{initial}}$-th percentile of sampled nonzero absolute objective differences and $p_{\text{initial}}$ is a given acceptance probability. Similarly, the final temperature is estimated as $T_{\text{low}}=\delta_{l}/\log(p_{\text{final}})$, where $\delta_{l}$ is the $\gamma_{\text{final}}$-th percentile of the same distribution and $p_{\text{final}}$ is a given acceptance probability. These energy differences are obtained by evaluating configurations drawn randomly from the solution space. We then form the temperature schedule following an exponential cooling scheme, where the temperature at the $i$-th annealing step is defined as $T_{i}=T_{\text{high}}\left(\frac{T_{\text{low}}}{T_{\text{high}}}\right)^{\frac{i}{i_{\max}-1}}$. Given the temperature sequence $\{T_{i}\}_{i=1}^{i_{\max}}$ and a current solution $x_{c}$, a candidate solution $x_{i}\in\mathcal{N}(x_{c})$ is accepted at iteration $k$ with probability $\min\left\{1,\ \exp\left[-\frac{(f(x_{c})+\Theta_{i})-f(x_{c})}{T_{k}}\right]\right\}$.
The parameters used for the heuristics is as follows. For the WalkQAP heuristic, we set the probability parameter $p=0.95$. For the Tabu heuristic, we set the maximum length of the tabu list to $20$, with the oldest entry being removed when a new neighbour is added whenever the list exceeds this limit. For the SA heuristic, objective value differences are sampled from 10 random configurations, each with 10 neighbours. The parameters settings for $T_{\text{high}}$ and $T_{\text{low}}$ are as follows: $p_{\text{initial}}=0.8$, $p_{\text{final}}=0.1$, $\gamma_{\text{initial}}=50$, and $\gamma_{\text{final}}=5$.
@@ 5.2 Comparison of Exact and Approximate Gradients
In this section, we discuss the results of our experiments using the heuristics listed in Section 5.1 for two settings of $\Theta$. The number of repeats was set to $100$, that is, $|\mathcal{T}|=100$. The Intel Xeon E7-8890 v3 running at 2.50 GHz with 144 cores was used for our experiments. Our aim is to assess the impact of omitting the error correction term $\mathcal{E}$ presented in Proposition 4.4. The primary motivation for omitting this step stems from its computational cost. Computing $\mathcal{E}$ involves a matrix–matrix multiplication $S_{34}Q$, which requires more-substantial processing resources than a matrix–vector multiplication. These operations significantly increase the overall computation time, especially as the problem size grows. The second reason is that when it comes to choosing a neighbour in a given full neighbourhood, the relative ordering of the values in $\Theta$ is important, not the values themselves. For example, given $\Theta^{1}=(-70,-40,-100,3,5)$ and $\Theta^{2}=(-80,-50,-90,3,5)$ representing two distinct gradients, we select the neighbour associated with the smallest value, which is the third element of each vector. Finally, IMC hardware devices typically have limited bit-precision capabilities. As a result, quantization of the nonzero elements in $Q$ becomes necessary. This introduces another potential source of error due to the loss of precision. Consequently, it is crucial to design algorithms that are resilient to multiple sources of error. Despite these challenges, incorporating the error correction step may still be feasible in certain scenarios, particularly when matrix–matrix multiplications can be efficiently performed.
Table 5.1 shows the minimum relative optimality gaps, $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{5})$, achieved by each heuristic under both approximate and exact gradient methods across 20 QAPLIB instances. Each value reflects the best relative gap obtained over $10^{5}$ iterations over 100 trials. For most instances, the Top10 and WalkQAP heuristics consistently achieve near-optimal values. However, Greedy and Tabu heuristics exhibit higher gaps when the approximate gradient is used. We analyze the behaviour of the Greedy, Tabu, and Top10 heuristics below.
Approximate Gradient Exact Gradient
Greedy SA Tabu Top10 WalkQAP Greedy SA Tabu Top10 WalkQAP
chr12a 97.30 0 97.30 0 0 3.81 0 0 0 0
chr15a 199.13 0 199.13 0 0 0 0 0 0 0
chr25a 184.88 2.32 184.88 0 0 16.33 0 0 0 0
esc16a 0 0 2.94 0 0 0 0 0 0 0
esc32e 200.00 0 200.00 0 0 0 0 0 0 0
had12 1.82 0.24 1.82 0 0 0 0 0 0 0
had20 1.59 1.47 1.53 0 0 0 0 0 0 0
lipa20a 3.58 3.34 3.58 0 0 1.06 0 0 0 0
lipa40a 2.03 2.57 2.03 0 0.74 1.25 0 0 0 0
nug20 4.90 2.49 4.75 0 0 0.54 0 0 0 0
nug30 6.30 1.53 5.98 0 0 1.01 0 0 0 0
sko42 4.45 2.57 3.45 0 0 1.39 0 0 0 0
sko49 4.06 1.98 4.20 0 0.13 1.17 0 0 0.01 0.06
tai12a 2.08 0 2.08 0 0 0 0 0 0 0
tai25a 4.07 5.77 4.07 0.89 0 1.91 0 0 0 0.41
tai40a 4.28 6.79 4.28 0.81 1.34 2.16 0.33 0.12 0.33 0.84
tai50a 4.37 7.16 4.37 0.94 1.18 2.91 0.75 0.50 0.72 1.34
tho30 3.36 1.41 3.36 0 0 1.39 0 0 0 0
tho40 6.02 3.06 6.02 0 0.05 2.19 0 0 0 0
wil50 1.69 2.20 1.10 0 0.02 0.62 0 0 0 0.02
Table 5.1: Minimum relative optimality gaps observed over $10^{5}$ iterations using the five heuristics.
5.2.1 Typical Behaviour of the Greedy and Tabu Search Heuristics
Given a feasible binary solution $x$, the Greedy heuristic iteratively selects the neighbour within $\mathcal{N}(x)$ that yields the maximum immediate reduction in the objective function. While this strategy ensures the best possible improvement at each step within the incumbent full neighbourhood, if often leads to an oscillatory pattern, where the search alternates between two permutation assignments. As a result, the algorithm fails to make further progress towards a global optimal solution. This pattern is evident for five of the selected QAPLIB instances, as shown in Figure 5.1. We make observations for both settings of $\Theta$, that is, $\Delta$ and $\tilde{\Delta}$.
Figure 5.1: Objective value trajectories of the Greedy heuristic on five QAPLIB instances over 50 iterations of full-neighbourhood evaluations, beginning with the identity permutation $X=I$.
From the sixth iteration of the QAPLIB instance “chr12a”, when the exact gradient is used, the Greedy search heuristic alternates between the permutation assignments $[2,6,0,1,4,5,7,11,3,9,8,10]\in\Pi_{12}$ and $[2,6,0,1,4,5,7,11,3,9,10,8]\in\Pi_{12}$, yielding the objective values 17,210 and 17,278, respectively. Similarly, when the approximate gradient is used, the Greedy heuristic alternates between $[0,5,2,3,4,1,6,7,8,9,10,11]\in\Pi_{12}$ and $[1,5,2,3,4,0,6,7,8,9,10,11]\in\Pi_{12}$ from the first iteration, yielding the objective values $46,646$ and $43,420$, respectively. Such alternating behaviour is a commonly observed pattern, which indicates that the Greedy search heuristic cannot escape local minima.
Given a feasible binary solution $x$ and its gradient $\Theta$ evaluated at $x$, the Tabu heuristic begins by sorting the elements of $\Theta$. It then iteratively checks the condition $\Theta_{i}<f_{\text{best}}-\langle x,Qx\rangle$, where $\Theta_{i}$ corresponds to the $i$-th neighbour in $\mathcal{N}(x)$. If this condition is satisfied, then the heuristic accepts that neighbour and proceeds to the next iteration. If $\Theta=\Delta$, then the condition reflects that the correct decision has been made, as $\Delta_{i}$ represents the true difference in the objective values. However, when an approximate gradient $\Theta=\tilde{\Delta}$ is used, discrepancies can occur due to gradient estimation errors. Specifically, if the error term is omitted from $\tilde{\Delta}_{i}+\mathcal{E}_{i}<f_{\text{best}}-\langle x,Qx\rangle$, the comparison may lead to the acceptance of an incorrect neighbour. In other words, decisions based solely on $\tilde{\Delta}_{i}$ may be unreliable in the presence of approximation errors. This behaviour is illustrated in Figure 5.2. When the exact gradient is used, the Tabu heuristic avoids cycling between configurations. In contrast, with the approximate gradient, its behaviour resembles that of the Greedy heuristic for the five instances shown in the figure, alternating between suboptimal configurations.
Figure 5.2: Objective value trajectories of the Tabu heuristic for five QAPLIB instances over 50 iterations of full-neighbourhood evaluations, beginning with the identity permutation $X=I$.
Consequently, we do not consider the Greedy and Tabu heuristics for further analysis.
5.2.2 Analysis of the Top10 Heuristic
We now describe the behaviour of the Top10 heuristic. Figure 5.3(a) shows a comparison of the relative optimality gaps using the approximate gradient $\tilde{\Delta}$ and exact gradient $\Delta$ observed over $10^{5}$ iterations, that is, $\left\{{\operatorname*{\text{rel\_gap}}}^{t}(10^{5}):t\in\mathcal{T}\right\}$. Figure 5.3(b) shows the best observed optimality gaps, $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{4})$ and $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{5})$, for both $\tilde{\Delta}$ and $\Delta$.
@@ 5.2.2 Analysis of the Top10 Heuristic
Both approaches yield optimality gaps that are smaller than $1\%$, demonstrating their effectiveness in reaching optimal or near-optimal solutions. This observation supports the use of the approximate gradient as a viable alternative to the exact gradient.
Figure 5.4 summarizes the number of times the approximate gradient gradient method outperforms the exact gradient method, and vice versa, across 100 independent trials for each instance.
Figure 5.4: Proportion of 100 trials in which the approximate or exact gradient method reaches the optimal value in fewer iterations when the Top10 heuristic is used.
We record how often each method reaches the optimal value. The relative performance varies by instance, indicating that no single method is universally superior. The results highlight that using the approximate gradient can be effective, and, in some cases, competitive with the exact method. Appendix B provides a comprehensive set of additional results where the approximate gradient is used for 117 QAPLIB instances.
However, when comparing the relative runtime of computing the exact versus approximate gradient methods, the advantage of using the approximate gradient becomes evident. Figure 5.5 shows the relative runtime, computed as the ratio of the time taken using $\Delta$ divided by the time taken using $\tilde{\Delta}$ for each instance, with the instances sorted by size. The plot shows that using the exact gradient method requires significantly more computational resources and takes considerably longer than the approximate gradient method. Notably, the relative time increase becomes more pronounced as the instance size increases.
Figure 5.5: Ratio of the average runtime using the exact gradient method versus the approximate gradient method across $100$ trials.
5.2.3 Assessment of Omission of the Error Correction Terms
@@
Figure 5.6 shows rel_error over the first $1000$ iterations for instances formulated by Skorin–Kapov (“sko”), Li and Pardalos (“lipa”), and Taillard (“tai”). We observe a consistent trend across all three classes, indicating that the contribution of $\mathcal{E}_{i}$ relative to $\Delta_{i}$ diminishes as the problem size increases, hence, the impact of omitting the error terms becomes negligible in large-scale instances.